If you invest $19,500 today, how much will you have in each of the following instances? Use Appendix A as an approximate answer, but calculate your final answer using the formula and financial calculator methods.


a. In 12 years at 11 percent? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  


b. In 18 years at 7 percent? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  


c. In 25 years at 8 percent? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  


d. In 10 years at 6 percent (compounded semiannually)? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
Answer #1

a.Information provided:

Present value= $19,500

Time= 12 years

Interest rate= 11%

The question is calculated by computing the future value.

Enter the below in a financial calculator to compute the future value:

PV= -19,500

N= 12

I/Y= 11

Press the CPT key and FV to compute the future value.

The value obtained is 68,219.77.

Therefore, I will have $68,219.77 in 12 years.

b. Information provided:

Present value= $19,500

Time= 18 years

Interest rate= 7%

The question is calculated by computing the future value.

Enter the below in a financial calculator to compute the future value:

PV= -19,500

N= 18

I/Y= 7

Press the CPT key and FV to compute the future value.

The value obtained is 65,908.68.

Therefore, I will have $65,908.68 in 18 years.

c. Information provided:

Present value= $19,500

Time= 25 years

Interest rate= 8%

The question is calculated by computing the future value.

Enter the below in a financial calculator to compute the future value:

PV= -19,500

N= 25

I/Y= 8

Press the CPT key and FV to compute the future value.

The value obtained is 133,545.27

Therefore, I will have $133,545.27 in 25 years.

d. Information provided:

Present value= $19,500

Time= 10 years*2= 20 semi-annual periods

Interest rate= 6%/2= 3% per semi-annual period

The question is calculated by computing the future value.

Enter the below in a financial calculator to compute the future value:

PV= -19,500

N= 20

I/Y= 3

Press the CPT key and FV to compute the future value.

The value obtained is 35,219.17.

Therefore, I will have $35,219.17 in 10 years.

In case of any query, kindly comment on the solution